OpenStudy (anonymous):

Find the exact value of the given expression using a sum or difference formula. 1) sin345° 2) cos(11π)/12

OpenStudy (anonymous):

1) sin(345) = sin(300+45)=sin(300)cos(45)+cos(300)sin(45)=(-sqrt(3)*sqrt(2))/(2*2) + (1*sqrt(2))/(2*2)=(-sqrt(6) + sqrt(2))/4

OpenStudy (anonymous):

2) cos(11pi/12)=cos (2π/3 + π/4) = cos 2π/3 cos π/4 - sin 2π/3 sin π/4 =-½*(√(2)/2) - √(3)/2* (√(2)/2) = (-sqrt(2)-sqrt(6))/4

OpenStudy (anonymous):

OK so I'm a little confused because my answer choices for #1 are: a.(√(3)+1)/(2√(2)) b.(√(3)-1)/(2√(2)) c.(-√(3)+1)/(2√(2)) d.(-√(3)-1)/(2√(2))

OpenStudy (anonymous):

Just factor out a sqrt(2) from my answer, and you should be able to find the right choice. :D

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

so it would be c then right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sweet! Gracias Senor!

OpenStudy (anonymous):

No problem. I'd appreciate it if you fanned me. :D

OpenStudy (anonymous):

Sure can do! Now your answer with three out of the same choice of answers doesn't match up though.. I don't think....

OpenStudy (anonymous):

I mean with number 2....

OpenStudy (anonymous):

What are your choices?

OpenStudy (anonymous):

same as the choices for #1...

OpenStudy (anonymous):

Factor out a sqrt(2). The answer would be d.

OpenStudy (anonymous):

Ohhhh I gotcha now lol sorry about that

OpenStudy (anonymous):

Haha np.